A solution contains a mixture of pentane and hexane at room temperature. The solution has a vapor pressure of 240 torr . Pure pentane and hexane have vapor pressures of 425 torr and 151 torr, respectively, at room temperature.
What is the mole fraction of hexane? (Assume ideal behavior.)
1
Expert's answer
2020-03-05T07:57:17-0500

"vapour\ pressure =P^o_{hex}*\chi_{hex}+P^o_{pen}*\chi_{pen}"

"Vapour \ pressure = 151*\chi_{hex}+425*\chi_{pen}"

"240=151*\chi_{hep}+425*(1-\chi_{hep})"

"(425-151)\chi_{hep}=425-240
\newline
\chi_{hep}=185/274= 0.675"